OpenStudy (anonymous):

Calculate the derivative of y=4log(base3)(x^2-1)

OpenStudy (anonymous):

Differentiate \[y=4\log_{3} (x ^{2}-1)\]

OpenStudy (anonymous):

since \[\log_b(x)=\frac{\ln(x)}{\ln(b)}\] by the change of base formula, you have \[\frac{d}{dx}\log_b(x)=\frac{1}{x\ln(b)}\]

OpenStudy (anonymous):

So it would be \[\frac{ 1 }{ (x ^{2} -1)\ln 3}\]

OpenStudy (anonymous):

@satellite73